Billy's father is twice billy's age. Eleven years ago, Billy's father was 3 times Billy's age. How old is Billy: Billy's father is twice billy's age. Eleven years ago, Billy's father was 3 times Billy's age. How old is Billy
Answer by munisa(5)

Let the present age of Billy be 'x' and the present age of his father 'y'.
Since Billy's father is twice as old as Billy,so
y=2x
11 years ago:
Billy's age was x-11 years.
Billy's father's age was y-11 years.
Since father was thrice as old,therefore:
y-11=3*(x-11)
Substituting for y we get,
2x-11=3*(x-11)
2x-11=3x-33
x=22
Hence, Billy is 22 years old.